A yardstick measures 0.5 by 2 by 36 inches. How many yardsticks will fit in a box 2 inches wide and 36 inches high, if the length of the box is 24 inches?
Answer by checkley77(12844)

2*24=48 in^2 is the area to be filled.
.5*2=1 in^2 is the area of each yard stick.
48/1=48 yard sticks can be placed in this box.